5100 invested for 10 years at 10 perceent compounded annually will accumulate to?
Answer by stanbon(75887)

5100 invested for 10 years at 10 percent
compounded annually will accumulate to?
--------------------
A(t) = P(1+(r/n))^(nt)
A(10) = 5100(1+(0.10/1))^(1*10)
----
A(10) = 5100(1.1)^10
A(10) = 5100*2.5937
---
A(10) = $13,228.09
-----------------------
Cheers,
Stan H.